Solve the following initial-value problem.

dp/dt = -kp, p(0) = 10.4, p(2) = 3.7


1
Expert's answer
2021-09-10T08:44:14-0400
"dp\/dt=-kp"

"\dfrac{dp}{p}=-kdt"

Integrate


"\int\dfrac{dp}{p}=-\int kdt"

"\ln|p|=-kt+\ln C"

"p=Ce^{-kt}"

The initial conditions


"p(0)=10.4=>10.4=C"

"p=10.4e^{-kt}"

"p(2)=3.7=>3.7=10.4e^{-k(2)}"

"k=0.5\ln(\dfrac{10.4}{3.7})"

"p=10.4(\dfrac{10.4}{3.7})^{-0.5t}"